The following trial balance of Concord Co. does not balance. CONCORD CO. TRIAL BALANCE JUNE 30, 2020 Debit Credit Cash $3,035Prepare a correct trial balance. CONCORD CO. Trial Balance Debit Credit D Totals $
Answer #1

Solution:

CONCORD CO

Trial Balance

June 30,2020

   Debit Credit
Cash 3,055
Accounts receivable 2,916
Supplies 300   
Equipment 4,630
Accounts payable 2,365
unearned service revenue 875
Common stock 6,165
Dividends 740
Retained earnings 3,165
Service revenue 3,836
Wages expense 3,660
Office expense 1,105
Total 16,406 16,406

Explanation:

Cash = 3035 + 750 - 570 - 80 - 80 = 3,055

Accounts receivable = 3096 - 750 + 570 =2,916

Supplies = 965-665 = 300

Equpment = 3965 + 665 = 4,630

Accounts payable = 2831 - 206 - 260 = 2365

unearned service revenue = 1365 - 490 = 875

Service revenue = 2545 + 890 - 89 + 490 = 3,836

Wages expense = 3565 + 835 - 740 = 3660